OpenStudy (anonymous):

HELPP !!! Write the equation of a quadratic whose roots are -3/7 and 2/5. (answer without fraction)

OpenStudy (anonymous):

\[-\frac{ 3 }{ 7 } and \frac{ 2 }{ 5 }\]

OpenStudy (jdoe0001):

\(\large {\begin{array}{llll} x =-\cfrac{3}{7}\qquad and \qquad &x=\cfrac{2}{5}\\ \quad \\ x +\cfrac{3}{7}=0&x-\cfrac{2}{5}=0\\ \quad \\ \left(x +\cfrac{3}{7}\right)=0&\left(x-\cfrac{2}{5}\right)=0 \end{array} \\ \quad \\ \textit{multiply both roots to get your polynomial}\\ \quad \\ \left(x +\cfrac{3}{7}\right)\left(x-\cfrac{2}{5}\right)=0}\)